# LEGO_HQEC: A Software Tool for Analyzing Holographic Quantum Codes

### v1

@@ II.2 Holographic Quantum Codes
Holographic quantum codes are a relatively new class of concatenated subsystem codes [62, 52, 63]. These codes have been found to exhibit properties consistent with the famous AdS/CFT correspondence [41, 43, 42, 64], in addition to possessing interesting and potentially useful error-correction properties for practical quantum computing, such as: high, tunable rates [52, 48]; versatility in constructions involving several different seed codes [62]; distance scaling better than $\sqrt{n}$ [54]; and constructions with large fault-tolerant logical gate sets [62], among other useful properties [55, 65].
Figure 3: Examples of several holographic tensor-network codes that LEGO_HQEC can construct and test. (a) The maximum-rate Harlow-Preskill-Pastawski-Yoshida (HaPPY) code on a hyperbolic pentagon tiling, mapping from logical (red) to physical (black) legs (Section IV.1.1). (b) A partially gauge-fixed version of the HaPPY code is created by projecting some logical legs onto an eigenstate |$\ket{G}$ of a gauge operator $G$. (c) The asymptotically zero-rate holographic Steane code, which uses the same 8-leg tensor once as a logical-to-physical map (central heptagon tensor) and otherwise with only planar legs (octagon tensors); a test of this code is performed in Section IV.1.2.
@@
A holographic quantum code is built typically from either one or several seed codes, as shown in Figure 3. As hyperbolic tessellations allow for many different orientations and polygonal embedments in the bulk according to the famous Schläfli symbol formula $\frac{1}{p}+\frac{1}{q}<\frac{1}{2}$, infinitely many variations of holographic quantum codes are possible to construct [62, 51]. However, only homogeneous [53, 40, 49, 52] (that is, those holographic codes constructed from uniform hyperbolic tessellations with only one polygon featured, as shown in Figure 3(a)) and simple heterogeneous (with only alternating layers of two seed codes, as in Figure 3(b)) holographic codes [62] have so far been constructed. Growing a holographic code layerwise is done in two distinct methods, as shown in Figure 4: either using (a) edge inflation, in which every edge is surrounded by vertices in the forthcoming layers of the tensor network, or (b) vertex inflation, in which each vertex is surrounded by edges in each subsequent layer. In the tensor-network scheme, the concatenation is apparent since subsequent layers must contract indices of the previous layer in order to adhere to the stringent constraints of a hyperbolic tessellation. A diagrammatic representation of these two methods can be seen in Figure 4. More details on this topic and in general on holographic codes can be found in [13].
@@
Figure 4: Edge inflation (a) and vertex inflation (b) of a tensor network on the $\{5,4\}$ face-centered hyperbolic tessellation [44, 45, 13]. In both cases, we show the first three inflation steps around a central pentagon.
@@ III Software Design
LEGO_HQEC was developed in Python, with ease of use front and center in the design considerations. The overarching design premise of LEGO_HQEC involves automation and extendability. As such, LEGO_HQEC is modularized into two parts: (a) the stabilizer and logical operator generator, and (b) the decoders. Both will be addressed in what follows.
III.1 Stabilizer & Logical Operator Generation
Stabilizer and logical operator generation is carried out following several steps, which are depicted in Figure 5. Firstly, the object TensorNetwork is defined, with dependencies Tensor and TensorLeg as important classes; these are the components used to construct the holographic tensor network. Next, the Push function is defined; in this portion, each tensor has its own OperatorPush method on its local legs. Push then activates tensors one-by-one from the center to the boundary, and sequentially generates boundary representations of all locally-defined stabilizers and logical operators. Additionally, legs that have undergone an operator push will be blocked, in order to prevent operators from being accidentally pushed back in towards the center; in this way a locally-defined operator can be pushed to the boundary of the network. Finally, the ReadOut layer retrieves boundary stabilizers and logical operators using a backtracking method, and compiles them into the final boundary result; this backtracking method starts from the center of the tensor network, then uses a depth-first traversal to reach outer tensors, while recording operators on the legs of the outermost tensors at the boundary. With this method, it is ensured that boundary operators align with geometric intuition from holographic codes. A a last step, a correctness check is performed by again traversing all internal legs and checking to see that all bulk-defined operators have been pushed to the boundary. More details on this portion of the software are available at [65].
@@ III.1 Stabilizer & Logical Operator Generation
III.2 Decoders
Three decoders are included in LEGO_HQEC: an erasure decoder (Section III.2.1) for the quantum erasure channel [8, 9], as well as two decoders for Pauli noise based on integer optimization and a tensor-network decoder (Section III.2.2) [48, 49].
III.2.1 Quantum Erasure Channel
Erasure (or qubit loss) is a quantum noise channel $\mathcal{E}_{e}(\cdot)$ with two inputs (|$\ket{0},\ket{1}$) and three outputs (|$\ket{0},\ket{1},\ket{e}$) [8, 66, 9], and can be written as
$\mathcal{E}_{e}(\rho)=(1-\epsilon)\rho+\epsilon\text{Tr}\big{[}\rho\big{]}\ket{e}\bra{e}~{},$ (21)
where with some probability $(1-\epsilon)$ a qubit is left unchanged and with probability $\epsilon$ the qubit is “erased”. |$\ket{e}$ is a state orthogonal to the input state $\rho$.
LEGO_HQEC includes an erasure decoder based on Gaussian elimination (GE), as was utilized in [52, 53]. The general idea of the GE decoder involves stacking all of the stabilizers and logical operators into a large matrix, and then using the support of the erasure vector, for which entries are randomly generated per qubit with probability $p$, to filter out columns that have undergone an erasure. The algorithm then proceeds by executing Gaussian elimination to find the reduced-row-echelon form of the matrix; if it is found that both logical operators still have physical support on the remaining qubits, then the trial is counted as a success. The runtime complexity of the GE algorithm is known to scale as $\mathcal{O}(n(n-k)^{2})$. More details on the GE decoder can be found at [53, 52].
III.2.2 Pauli Noise Channels
The prototypical Pauli noise channel is represented by the depolarizing noise channel $\mathcal{E}_{d}(\cdot)$ [9, 8], and takes the form
@@ III.2.2 Pauli Noise Channels
where $\delta$ is the probability of an error occurring, and $\{X,Y,Z\}$ represent the three Pauli operators. For Pauli noise, LEGO_HQEC employs two different types of decoders: one based on integer optimization [52, 54], and another based on tensor-network decoding [55, 49].
Integer optimization decoding works by a very similar technique as for the GE decoder discussed in Section III.2.1, but with some small changes. Firstly, as error locations are not known, one must first calculate the Moore-Penrose pseudoinverse [67, 68], which corresponds to the destabilizers (or pure errors) of the code [54, 69]. One then obtains a guessed error by multiplying the pseudoinverse matrix directly by the vector of the obtained syndrome. Finally, an integer optimization is utilized in order to deduce the lowest-weight pure-error correction [65, 54]. It is known that integer-optimization decoding exhibits with a runtime complexity of $\mathcal{O}(2^{n})$, which is dependent on the size of the quantum code. More details on integer-optimization decoding can be found in [65, 52, 54].
Tensor-network decoding, however, works by exploiting tensor-network geometry and clever contraction sequences [49, 48]. By setting the bond dimension of each tensor in the network to be four, Pauli errors $\{I,X,Y,Z\}$ can be easily accounted for. In this way, an error vector is initialized at random with some error probability $d$. These error probabilities are then utilized in the tensor network, and, upon contraction, yield a close approximation in the thermodynamic limit to maximum-likelihood decoding [50]. It is known that tensor-network decoding exhibits a runtime complexity of $\mathcal{O}(n^{2.37})$ in the best-case scenario [70, 71], and that multiple logical qubits can be decoded in parallel as $\mathcal{O}(nk)$, as long as the error rate stays below the critical error rate of the central logical qubit (i.e., the threshold of the central logical qubit) [48]. More details on tensor-network decoding can be found in [47, 49, 48].
@@ IV Examples Using LEGO_HQEC
This section contains a step-by-step set of examples for utilizing LEGO_HQEC. We begin by initializing and testing a maximum-rate Harlow-Preskill-Pastawski-Yoshida (HaPPY) [40] code under erasure noise, and then following up with an asymptotically zero-rate Steane code [53] under Pauli depolarizing noise. The results of both decoder simulations are shown in Figure 6.
@@
Figure 6: (a) Erasure decoder results for the maximum-rate HaPPY code (which obtains no threshold, as per the results in [40]), which utilized the GE decoder (Section III.2.1). (b) Pauli depolarizing noise results from the tensor-network decoder for an asymptotically zero-rate Steane code; the threshold of this code is found to be approximately $18.25\%$, in agreement with results from [54]. $750$ and $500$ Monte Carlo trials were utilized for the erasure and depolarizing noise simulations, respectively. Results are gathered per layer of concatenation $L$.
IV.1 Usage
IV.1.1 Maximum-Rate HaPPY Code under Erasure Noise
1from OperatorPush.PushingToolbox import batch_push, batch_push_multiprocessing
2from QuDec.InputProcessor import extract_stabilizers_from_result_dict, extract_logicals_from_result_dict
3from QuDec.OutputProcessor import save_results_to_csv
4import numpy as np
@@ IV.1.1 Maximum-Rate HaPPY Code under Erasure Noise
5from QuDec.ErasureDecoder import calculate_recovery_rates_for_p_range
@@
6from OperatorPush.Presets.Zero_Rate_HaPPY_new_for_kj import setup_zero_rate_happy, setup_max_rate_happy
7if __name__ == ’__main__’:
8 # HaPPY Code, max rate
9 for R in [0, 1, 2]:
10 tensor_list = setup_max_rate_happy(R=R)
11 results_dict = batch_push_multiprocessing(tensor_list)
12 stabilizers = extract_stabilizers_from_result_dict(results_dict)
13 logical_zs, logical_xs = extract_logicals_from_result_dict(results_dict)
14 print("logical_xs: ", logical_xs, logical_zs)
15 logical_operators = [logical_zs[0]] + [logical_xs[0]]
16 results = calculate_recovery_rates_for_p_range(n=750, p_start=0.01, p_end=1.0, p_step=0.01, stabilizers=stabilizers,logical_operators=logical_operators,n_process=1, cpu_affinity_list=None)
17 save_results_to_csv(results, file_path=F’R{R}_rec_happy.csv’)
Listing 1: Code snippet for initializing the stabilizers and logical operators of the maximum-rate HaPPY code, as well as for testing the first three layers of concatenation under the erasure noise channel. Decoding was carried out using the GE decoder.
LABEL:code1 displays a code snippet for obtaining the threshold of the maximum-rate HaPPY code under erasure noise, utilizing the GE decoder discussed in Section III.2. In the example above, the commensurate libraries and dependencies are first imported (lines 1-6), and then the stabilizers and logical operators for each layer are generated (lines 10-13). Next, those objects are then fed into the erasure decoder per layer $L$, and the Monte Carlo trials commence layerwise (lines 9 and 16). Finally, the results are written into a csv file and saved (lines 19).
IV.1.2 Asymptotically Zero-Rate Steane Code under Depolarizing Noise
1from QuDec.TN_decoder import tn_quantum_error_correction_decoder_multiprocess
2import numpy as np
@@ IV.1.2 Asymptotically Zero-Rate Steane Code under Depolarizing Noise
3from OperatorPush.PushingToolbox import batch_push, batch_push_multiprocessing
@@
4from QuDec.InputProcessor import extract_stabilizers_from_result_dict, extract_logicals_from_result_dict
@@
5from QuDec.OutputProcessor import save_results_to_csv
6from OperatorPush.Presets.Heptagon_Steane_Code import setup_heptagon_zero_rate_steane
7if __name__ == ’__main__’:
@@
8 # Steane Code, zero rate
9 for R in [0, 1, 2]:
10 p_depo_step = 0.01
11 p_depo_start = 0.1
12 p_depo_end = 0.5
13
14 tensor_list = setup_heptagon_zero_rate_steane(R=R)
15 results_dict = batch_push_multiprocessing(tensor_list)
16 stabilizers = extract_stabilizers_from_result_dict(results_dict)
17 logical_zs, logical_xs = extract_logicals_from_result_dict(results_dict)
18 success_rates = []
19 rx = 0.33
20 rz = 0.33
21 ry = 1 - rx - rz
22 for p in np.arange(p_depo_start, p_depo_end, p_depo_step):
@@
23 results = tn_quantum_error_correction_decoder_multiprocess(tensor_list=tensor_list, p=p, N=500, rx=rx, ry=ry, rz=rz, stabilizers=stabilizers, logical_z=logical_zs[0], logical_x=logical_xs[0], n_process=1)
24 success_rates.append((p, results))
@@
LABEL:code2 shows the code needed for initializing the tensor-network decoder for a depolarizing noise channel. In contrast to the previous case, the tensor network decoder requires a specification of the relative proportions of X, Y, and Z errors (lines 19-21). However, the actual generation of the stabilizers and logical operators proceeds identically as for the HaPPY code (lines 14-17). In the future, functionality will be added in order to allow the user to specify which tensors will have usable logical qubits, as well as gauge-fixing patterns that have featured in other holographic quantum codes recently [52, 63].
IV.2 Runtime Performance
Here we report an example of the runtime performance expected of an asymptotically zero-rate HaPPY code for: the stabilizer and logical operator generation OperatorPush; the GE erasure decoder; the integer-optimization decoder; and for the tensor-network decoder. All results are given as a function of layer $L$ and times are reported in seconds (s). For the decoders, the plotted results are per Monte Carlo trial, while the OperatorPush results denote the per layer start-to-finish time for pushing all locally-defined stabilizers and logical operators to their respective boundary representations, and extracting them into a csv file.

### v3

@@ II-B Holographic Quantum Codes
Holographic quantum codes are a relatively new class of generalized concatenated codes [heterogeneous_holo_qrm, steinberg2024far, steinberg_2023]. These codes have been found to exhibit properties consistent with the AdS/CFT correspondence [harlow2015bulk, witten1998anti, maldacena1999large, gubser1998gauge], in addition to possessing interesting and potentially useful error-correction properties for practical quantum computing, such as: high, tunable rates [steinberg2024far, farrelly_parallel_decoding]; versatility in constructions involving several different seed codes [heterogeneous_holo_qrm]; distance scaling better than $\sqrt{n}$ [harris_int_opt]; and constructions with large fault-tolerant logical gate sets [heterogeneous_holo_qrm], among other useful properties [fan2024overcoming, junyu_thesis].
[…]
@@
A holographic quantum code is built typically from either one or several seed codes, as shown in Figure 3. As hyperbolic tessellations allow for many different orientations and polygonal embeddings in the bulk subject to the constraint $\frac{1}{p}+\frac{1}{q}<\frac{1}{2}$, infinitely many variations of holographic quantum codes are possible to construct [heterogeneous_holo_qrm, ratcliffe1994foundations]. However, only homogeneous [harris_css_hqec, preskill_happy_code, farrelly_tn_codes, steinberg2024far] (that is, those holographic codes constructed from uniform hyperbolic tessellations with only one polygon featured, as shown in Figure 3(a)) and simple heterogeneous (with only alternating layers of two seed codes, as in Figure 3(b)) holographic codes [heterogeneous_holo_qrm] have so far been constructed. Growing a holographic code layerwise is done in two distinct methods, as shown in Figure 4: either using (a) edge inflation, in which every inflation step adds the polygons adjacent to the edges of the previous boundary, or (b) vertex inflation, in which each inflation step adds the polygons adjacent to the boundary vertices, resulting in a connected ring of polygons to be added. In the tensor-network scheme, subsequent layers must contract indices of the previous layer in order to adhere to the stringent constraints of a hyperbolic tessellation. More details on this topic and the motivation behind holographic codes can be found in [jahn_topical].
@@
Figure 4: Methods by which one can perform holographic concatenation. Both (a) edge inflation and (b) vertex inflation of a tensor network on the $\{5,4\}$ face-centered hyperbolic tessellation [conformal_quasicrystals_holography, jahn2022tensor_qcft, jahn_topical]. In both cases, we show the first three inflation steps around a central pentagon.
@@ III Software Design
LEGO_HQEC was developed in Python. For ease of use, the code is modularized into two parts: (a) the stabilizer and logical operator generator, and (b) the decoders. Both will be addressed in what follows.
III-A Stabilizer & Logical Operator Generation
Stabilizer and logical operator generation is carried out following several steps, which are depicted in Figure 5. Firstly, the object TensorNetwork is defined, with dependencies Tensor and TensorLeg as important classes; these are the components used to construct the holographic tensor network. Next, the Push function is defined; in this portion, each tensor has its own OperatorPush method on its local legs. Push then activates tensors one-by-one from the center to the boundary, and sequentially generates boundary representations of all locally-defined stabilizers and logical operators. Additionally, legs that have undergone an operator push will be blocked, in order to prevent operators from being accidentally pushed back in towards the center; in this way a locally-defined operator can be pushed to the boundary of the network. Finally, the ReadOut layer retrieves boundary stabilizers and logical operators using a backtracking method, and compiles them into the final boundary result; this backtracking method starts from the center of the tensor network, then uses a depth-first traversal to reach outer tensors, while recording operators on the legs of the outermost tensors at the boundary. With this method, it is ensured that boundary operators align with the geometric intuition expected from holographic codes. As a last step, a correctness check is performed by again traversing all internal legs and checking that all bulk-defined operators have been pushed to the boundary. More details on this portion of the software are available in [junyu_thesis].
@@ III-A Stabilizer & Logical Operator Generation
III-B Decoders
Three decoders are included in LEGO_HQEC: An erasure decoder (Section III-B1) for the quantum erasure channel [wilde2013quantum, nielsen2010quantum] as well as two decoders for Pauli noise based on integer optimization and a tensor-network decoder (Section III-B2[farrelly_parallel_decoding, farrelly_tn_codes].
III-B1 Quantum Erasure Channel
[…]
$|,\mathcal{E}_{e}(\rho)=(1-\epsilon)\rho+\epsilon\text{Tr}\big[\rho\big]\ket{e}\bra{e}$~, (10)
[…]
LEGO_HQEC includes an erasure decoder based on Gaussian elimination (GE), as was utilized in [steinberg2024far, harris_css_hqec]. The general idea of the GE decoder involves stacking all of the stabilizers and logical operators into a large matrix, and then using the support of the erasure vector, for which entries are randomly generated per qubit with probability $p$, to filter out columns that have undergone an erasure. The algorithm then proceeds by executing Gaussian elimination to find the reduced-row-echelon form of the matrix; if it is found that both logical operators still have physical support on the remaining qubits, then the trial is counted as a success. The runtime complexity of the GE algorithm is known to scale as $\mathcal{O}(n(n-k)^{2})$. More details on the GE decoder can be found at [harris_css_hqec, steinberg2024far].
III-B2 Pauli Noise Channels
The prototypical Pauli noise channel is represented by the depolarizing noise channel $\mathcal{E}_{d}(\cdot)$ [nielsen2010quantum, wilde2013quantum], and takes the form
@@ III-B2 Pauli Noise Channels
where $\delta$ is the probability of an error occurring, and $\{X,Y,Z\}$ represent the three Pauli operators. For Pauli noise, LEGO_HQEC employs two different types of decoders: one based on integer optimization [steinberg2024far, harris_int_opt], and another based on tensor-network decoding [fan2024overcoming, farrelly_tn_codes].
Integer optimization decoding works via a very similar technique to the GE decoder discussed in Section III-B1, but with some small changes: Firstly, as error locations are not known, one must first calculate the Moore-Penrose pseudoinverse [penrose1955generalized, moore1920reciprocal], which corresponds to the destabilizers (or pure errors) of the code [harris_int_opt, gottesman_improved_sim]. One then obtains a guessed error by multiplying the pseudoinverse matrix directly by the vector of the obtained syndrome. Finally, an integer optimization is utilized in order to deduce the lowest-weight pure-error correction [junyu_thesis, harris_int_opt]. It is known that integer-optimization decoding exhibits a runtime complexity of $\mathcal{O}(2^{n})$, which is dependent on the size of the quantum code. More details on integer-optimization decoding can be found in [junyu_thesis, steinberg2024far, harris_int_opt].
Tensor-network decoding, however, works by exploiting the tensor-network geometry and clever contraction sequences [farrelly_tn_codes, farrelly_parallel_decoding]. By setting the bond dimension of each tensor in the network to be four, Pauli errors $\{I,X,Y,Z\}$ can be easily accounted for. In this way, an error vector is initialized at random with some error probability $d$. These error probabilities are then utilized in the tensor network, and, upon contraction, yield a close approximation in the thermodynamic limit to maximum-likelihood decoding [chubb2021statistical]. It is known that tensor-network decoding exhibits a runtime complexity of $\mathcal{O}(n^{2.37})$ [le_gall_matrix_mult, bravyi_suchara] [^1] Although the runtime complexity above has been reported in several studies, we note that the overall runtime complexity for tensor-network decoding depends strongly on the tensor network’s geometry, as well as the contraction scheme (i.e., whether or not approximate contractions are used) and contraction ordering., and that multiple logical qubits can be decoded in parallel as $\mathcal{O}(nk)$, as long as the error rate stays below the critical error rate of the central logical qubit (i.e., the threshold of the central logical qubit) [farrelly_parallel_decoding]. More details on tensor-network decoding can be found in [joe_chubb, farrelly_tn_codes, farrelly_parallel_decoding].
@@ IV Examples Using LEGO_HQEC
This section contains a step-by-step set of examples for utilizing LEGO_HQEC as well as a runtime analysis. We begin by initializing and testing a maximum-rate Harlow-Preskill-Pastawski-Yoshida (HaPPY) [preskill_happy_code] code under erasure noise, and then following up with an asymptotically zero-rate Steane code [harris_css_hqec] under Pauli depolarizing noise. The results of both decoder simulations are shown in Figure 6.
@@
Figure 6: (a) Erasure decoder results for the maximum-rate HaPPY code (which obtains no threshold, as per the results in [preskill_happy_code]), which utilized the GE decoder (Section III-B1). (b) Pauli depolarizing noise results from the tensor-network decoder for an asymptotically zero-rate Steane code; the threshold of this code is found to be approximately $18.25\%$, in agreement with results from [harris_int_opt]. $1000$ Monte Carlo trials were utilized for the erasure and depolarizing noise simulations. Results are gathered per layer of concatenation $L$. The code snippets related to both (a) and (b) can be found in LABEL:code1 and LABEL:code2, respectively. Regarding computation time, for (a) the recovery rate curves for $L=1$ to $L=5$ took $14$ s, $15$ s, $18$ s, $536$ s, and $2204$ s respectively. In (b) the cases $L=0$ to $L=2$ took $15$ s, $194$ s, and $1497$ s, respectively (on an i9-14900K processor using 24 threads).
IV-A Usage
IV-A1 Maximum-Rate HaPPY Code under Erasure Noise
@@
from LEGO_HQEC.QuDec.ErasureDecoder import calculate_recovery_rates_for_p_range
@@
from LEGO_HQEC.OperatorPush.Presets.HaPPY_code import setup_max_rate_happy
@@
from LEGO_HQEC.OperatorPush.PushingToolbox import batch_push
@@
from LEGO_HQEC.QuDec.InputProcessor import extract_stabilizers_from_result_dict, extract_logicals_from_result_dict
@@
from LEGO_HQEC.QuDec.OutputProcessor import save_results_to_csv
if __name__ == ’__main__’:
# Examine the HaPPY code with radii R=0, 1, 2, and 3, respectively.
for R in [1, 2, 3, 4, 5]:
# Obtain TN of the zero-rate HaPPY code with corresponding radius via preset.
tensor_list = setup_max_rate_happy(R=R)
results_dict = batch_push(tensor_list)
stabilizers = extract_stabilizers_from_result_dict(results_dict)
logical_zs, logical_xs = extract_logicals_from_result_dict(results_dict)
logical_operators = [logical_zs[0]] + [logical_xs[0]]
# Starting from an erasure error rate p of 0.0 to 1.0 with a step size of 0.05,
results = calculate_recovery_rates_for_p_range(n=1000, p_start=0.0, p_end=1.0, p_step=0.05, stabilizers=stabilizers, logical_operators=logical_operators)
save_results_to_csv(results, file_path=F’R{R}_rec.csv’)
Listing 1: Code snippet for initializing the stabilizers and logical operators of the maximum-rate HaPPY code, as well as for testing the first through fifth layers of concatenation under the erasure noise channel. Decoding was carried out using the GE decoder.
LABEL:code1 displays a code snippet for the construction and execution of the threshold calculation for the maximum-rate HaPPY code under erasure noise, utilizing the GE decoder discussed in Section III-B. In the example above, the commensurate libraries and dependencies are first imported (lines 1-5), and then the stabilizers and logical operators for each layer are generated (lines 11-13). Those objects are then fed into the erasure decoder per layer $L$, and the Monte Carlo trials commence layerwise (line 17). Finally, the results are written into a CSV file and saved (line 19).
@@
IV-A2 Asymptotically Zero-Rate Steane Code under Depolarizing Noise
from LEGO_HQEC.QuDec.TN_decoder import tn_quantum_error_correction_decoder_multiprocess
from LEGO_HQEC.OperatorPush.Presets.Holographic_Steane_code import setup_heptagon_zero_rate_steane
from LEGO_HQEC.QuDec.OutputProcessor import save_results_to_csv
import numpy as np
if __name__ == ’__main__’:
rx = 1/3
rz = 1/3
ry = 1/3
for R in [0, 1, 2]:
tensor_list = setup_heptagon_zero_rate_steane(R=R)
p_depo_step = 0.05
p_depo_start = 0.0
p_depo_end = 0.5
succ_rates = []
for p_depo in np.arange(p_depo_start, p_depo_end + p_depo_step, p_depo_step):
print(f"Running at R = {R}, p_depo = {p_depo}")
@@ IV-A2 Asymptotically Zero-Rate Steane Code under Depolarizing Noise
succ_rate = tn_quantum_error_correction_decoder_multiprocess(tensor_list=tensor_list, p=p_depo, rx=rx, ry=ry, rz=rz, N=1000, n_process=8)
succ_rates.append((p_depo, succ_rate))
@@
LABEL:code2 shows the code needed for initializing the tensor-network decoder for a depolarizing noise channel. In contrast to the previous case, the tensor network decoder requires a specification of the relative proportions of $X$, $Y$, and $Z$ errors (lines 7-9). However, the actual generation of the stabilizers and logical operators proceeds identically as for the HaPPY code (lines 14-17).
IV-B Runtime Performance
Here we report an example of the runtime performance expected of an asymptotically zero-rate HaPPY code for: the stabilizer and logical operator generation OperatorPush; the GE erasure decoder; the integer-optimization decoder; and the tensor-network decoder. All results are given as a function of layer $L$ and times are reported in seconds (s). For the decoders, the plotted results are per Monte Carlo trial, while the OperatorPush results denote the per-layer start-to-finish time for pushing all locally-defined stabilizers and logical operators to their respective boundary representations, and extracting them into a CSV file.